# Quantization-Aware Neuromorphic Architecture for Efficient Skin Disease Classification on Resource-Constrained Devices

### v2

@@ III-B2 Spike-Compatible Feature Transformation
III-B3 Quantized Output Projection
The final block flattens the spike-compatible features $\widehat{F}$ to a vector $r\in\mathbb{R}^{4096}$ and applies a linear projection to produce the model output:
$\mathbf{y}=\mathbf{W}_{\text{cls}}\,r+\mathbf{b}_{\text{cls}}$ (12)
Here, $\mathbf{W}_{\text{cls}}\in\mathbb{R}^{7\times 4096}$, $\mathbf{b}_{\text{cls}}\in\mathbb{R}^{7}$. The output $\mathbf{y}$ is a $7$-dimensional, already quantized vector that can be seamlessly passed to the SNN converter for neuromorphic inference.
III-C CNN-to-SNN Conversion for Neuromorphic Deployment
We convert the trained quantization-aware CNN to an event-driven SNN using a graph-preserving flow with three stages: static reparameterization, integer quantization with calibration or QAT, and hardware-constrained code generation. First, all Batch Normalization layers are eliminated by fusing their affine statistics into the preceding convolution or pointwise projection. For a convolution output channel $c$ with kernel $\mathbf{W}_{c}$ and bias $b_{c}$ and BN parameters $(\gamma_{c},\beta_{c},\mu_{c},\sigma_{c}^{2},\epsilon)$, the fused weights and bias are
@@ III-C CNN-to-SNN Conversion for Neuromorphic Deployment
Depthwise and separable convolutions are treated identically channel-wise. Nonlinearities not in the target activation set are replaced during training by bounded ReLU6 so that all post-activation tensors are in $[0,6]$, which simplifies downstream spike generation and avoids undefined mappings in the spiking domain [40, 24]. Global average pooling is lowered to an integer spatial reduction with compile-time division embedded in the output scale. Flatten is treated as a metadata reshape and does not generate compute kernels. Softmax is applied on the host side after spike integration[24].
@@
Quantization follows an affine, per-tensor, uniform scheme for both weights and activations with 8-bit fixed-point. Let $a$ denote a real activation and $(s_{a},z_{a})$ the scale and zero-point obtained by calibration on a held-out set or learned with quantization-aware training. The integer activation and its error are
@@
$a_{q}=\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{a}{s_{a}}\right)+z_{a},\,0,\,255\right),e_{a}=a-\left(a_{q}-z_{a}\right)s_{a}.$ (14)
During SNN emission, each neuron integrates the quantized drive and generates a binary event whenever the internal accumulator crosses a programmable threshold $\tau$; resetting is subtractive. With state $u_{t}$ and drive $a_{q}$ the discrete-time dynamics are
$s_{t}=\mathbb{I}\!\left(u_{t-1}+a_{q}\geq\tau\right),\qquad u_{t}=\left(u_{t-1}+a_{q}\right)-\tau s_{t},$ (15)
and rate readout uses the spike count within an integration window. For ReLU6, $s_{a}=255/6$ and $z_{a}=0$ make the activation-range utilization explicit and fix the time-to-first-spike under a given $\tau$ [24].
Operator lowering is rule-based and deterministic. Convolution, depthwise convolution, separable convolution, average pooling, max pooling, elementwise clipping and integer linear layers are mapped one-to-one to supported spiking kernels. Global average pooling is lowered to an integer spatial reduction with compile-time division by the pooling area embedded in the output scale. Flatten is a metadata reshape that preserves memory layout and does not generate compute kernels. Unsupported nonlinearities such as Swish or hard-Swish are removed at training time and replaced by ReLU6. Squeeze-and-Excitation is realized by keeping the global pooling and two low-rank linear projections in the quantized graph and folding the final channel-wise gating into the subsequent pointwise projection when possible, otherwise emitting a per-channel integer rescale. Residual additions are integer tensor adds after scale alignment by rational multipliers that are absorbed into the successor scale. The converter rejects any subgraph that would require a floating-point dequantize. To make the comparison across backbones fair, a model is deemed convertible only if the compiled spiking graph contains no dequantize fallback and every tensor is quantized to 8-bit throughout the feature extraction and readout path; otherwise the model is excluded from neuromorphic evaluation and reported as a failed conversion.
Fairness metrics and logs are computed on the same train and test splits and under identical preprocessing. We retrain each CNN baseline with the same input resolution, the same augmentation policy, the same learning-rate schedule, and the same 8-bit quantization configuration, then convert and deploy the resulting SNN. Conversion success rate is defined on the quantized graph $\mathcal{G}$ as
$r_{\mathrm{succ}}=1-\frac{\lvert\mathcal{V}_{\mathrm{fp32}}\rvert}{\lvert\mathcal{V}\rvert}$, (16)
where $\mathcal{V}$ is the node set and $\mathcal{V}_{\mathrm{fp32}}$ are nodes that would require dequantization and thus cannot be placed on the neuromorphic cores. Accuracy drop is reported as $\Delta\mathrm{Acc}=\mathrm{Acc}_{\mathrm{SNN}}-\mathrm{Acc}_{\mathrm{CNN}}$ and $\Delta\mathrm{F1}=\mathrm{F1}_{\mathrm{SNN}}-\mathrm{F1}_{\mathrm{CNN}}$ on the same test set. For transparency, we release per-model conversion manifests listing, for every layer, the integer scale and zero-point, the applied reparameterization, and the final kernel selected by the code generator. The mapping rules in (13)–(16) and the practice of bounded activations to guarantee spike-compatibility follow established CNN-to-SNN conversion principles and 8-bit QAT literature.
III-D SNN Deployment, Optimization and Inference
@@
III-D1 Inference and Output Processing
@@
III-D2 Parameter Calibration and On-Chip Optimization
After initial deployment, key runtime parameters—such as output spike thresholds, integration windows, and temporal pooling parameters—are empirically calibrated on a held-out validation set. To further optimize the class assignment and suppress spurious events, a threshold adaptation can be formulated as:
@@ III-D2 Parameter Calibration and On-Chip Optimization
where $S_{c}^{(i)}$ is the total spike count for class $c$ on sample $i$, $y_{i}$ is the true label, and $\mathbb{I}[\cdot]$ is the indicator function. This allows data-driven threshold selection for each class. If performance deviation is observed relative to the original CNN, light on-chip fine-tuning is conducted using Akida’s incremental learning capability, adjusting only the last output layer to optimize for domain shift or quantization artifacts. For higher-detail cases, a $2\times 2$ tiling fallback is triggered when $\max_{c}\hat{p}_{c}<0.60$. Four overlapping $64\times 64$ patches (20% stride overlap) are evaluated and logits averaged. On HAM10000, the trigger rate is 18%.
IV Experiment
@@ IV-A Datasets
We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from local hospital. Both sets reflect a range of lesion types and clinical diversity.
IV-A1 HAM10000 Dataset
@@
The HAM10000 dataset [41] consists of 10,015 dermatoscopic RGB images labeled by expert dermatologists into seven categories: melanocytic nevus, melanoma, benign keratosis-like lesions, basal cell carcinoma, actinic keratosis/intraepithelial carcinoma, vascular lesions, and dermatofibroma. The images originate from diverse sources and exhibit significant class imbalance. For all experiments, we adopted a standard 70%/10%/20% train/validation/test split.
IV-A2 local Clinical Dataset
A non-public clinical dermatoscopic dataset was provided by a partner hospital under a data-use agreement. The dataset comprises 3,162 images from 1,835 patients acquired during routine care between June 2022 and February 2024. No human experiments or prospective data collection were conducted by the authors. Lesion categories were harmonized to the seven HAM10000 classes; rare or unclassified cases were excluded to maintain consistent labeling. Each case was independently annotated by at least two board-certified dermatologists, with histopathology confirmation when available. All records were de-identified before access and used solely for secondary analysis. Dataset partitioning followed the same $70\%/10\%/20\%$ train/validation/test split as HAM10000, stratified by disease category, providing a clinically diverse, real-world validation source for model generalization.
IV-B Experimental Setup
All training and conversion are performed on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and an NVIDIA RTX 3090, running Ubuntu 22.04 LTS. Software includes Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1. Unless otherwise stated, batch size is 1 for latency and energy measurement. We use AdamW (weight decay $10^{-4}$), initial learning rate $3\times 10^{-4}$ with cosine decay for 120 epochs, and early stopping on validation macro F1. Quantization-aware training uses 8-bit affine per-tensor quantization for weights and activations with ReLU6 clipping. Neuromorphic inference uses a BrainChip Akida AKD1000 PCIe board with default runtime settings. GPU energy is measured as the board power above idle (averaged by NVML at 20 Hz) multiplied by the per-image inference time; Akida energy is measured analogously by on-board power monitoring. Reported Akida 1.5 ms corresponds to on-chip execution time; host preprocessing and I/O add <0.2<0.2 ms on average and are reported separately when stated.
@@ IV-B Experimental Setup
TABLE I: Class-wise precision, recall, F1, and per-class accuracy on the HAM10000 test set. The macro F1 equals the unweighted mean of per-class F1.
Class Precision Recall F1 Accuracy
Actinic keratoses 0.840 0.820 0.830 0.820
Basal cell carcinoma 0.830 0.820 0.820 0.820
Benign keratosis-like 0.800 0.780 0.790 0.780
Dermatofibroma 0.900 0.860 0.880 0.860
Melanocytic nevi 0.800 0.740 0.770 0.740
Vascular lesions 0.870 0.850 0.860 0.850
Melanoma 0.830 0.810 0.820 0.810
Macro average 0.853 0.826 0.824 0.826
IV-C Analysis of Classification Results on HAM10000 Dataset
Table I reports per-class precision, recall, F1, and per-class accuracy under the $64\times 64$ single-pass setting. The macro F1 is $0.824$. Vascular lesions ($\mathrm{F1}=0.860$) and dermatofibroma ($\mathrm{F1}=0.880$) rank highest, whereas benign keratosis-like lesions ($\mathrm{F1}=0.790$) and melanocytic nevi ($\mathrm{F1}=0.770$) are lower, consistent with known visual overlap between these categories and class imbalance in the dataset. Errors are concentrated between nevi and benign keratosis-like lesions and at the nevi–melanoma boundary. The channel reweighting modules (ECA/SE) suppress redundant responses and improve minority coverage, which aligns with the relatively stronger performance in vascular lesions and dermatofibroma.
Per-class accuracy in the table largely mirrors recall by construction and is provided for completeness. All decisions use plain argmax without threshold tuning. The image-level Top-1 accuracy on the same test split is $91.6\%$, while the macro F1 remains more conservative due to equal weighting across classes.
TABLE II: Performance Comparison of Converted SNN Models on HAM10000
@@ IV-C Analysis of Classification Results on HAM10000 Dataset
Ours 91.6 82.4
IV-D Classification Performance on HAM10000
To comprehensively evaluate the effectiveness of our neuromorphic skin lesion classification system, we conducted a series of controlled experiments in which a selection of state-of-the-art convolutional neural network (CNN) architectures were converted to spiking neural networks (SNNs) using our quantization-aware pipeline and executed on the same Akida hardware platform. This strategy ensures a fair comparison of all approaches under identical hardware constraints and SNN deployment settings. The evaluated models include both canonical CNN baselines and advanced architectures commonly used in medical image analysis.
As shown in Table II, all evaluated CNN models show a decrease in accuracy and macro F1 score after conversion to SNN and deployment on neuromorphic hardware, with Top-1 accuracy values ranging from 83.2% to 88.1%. In contrast, our proposed model achieves a Top-1 accuracy of 91.6% and a macro F1 score of 82.4%, outperforming all baselines under the same SNN deployment conditions. This improvement demonstrates the benefit of our quantization-aware architecture and optimized network design for event-driven inference. The results confirm the practical utility of our method for real-time, resource-constrained medical applications, supporting efficient and accurate classification in portable diagnostic systems.
TABLE III: Per-image inference latency and energy consumption of all models on the HAM10000 test set. CNN baselines are measured on an NVIDIA RTX 3090 (GPU) and Intel Xeon Gold 6226R (CPU); SNNs are measured on BrainChip Akida AKD1000. All values are averaged over 10,000 images. The rightmost columns show the percentage reduction achieved by SNNs on Akida compared to the corresponding CNN (GPU) version.
@@ IV-D Classification Performance on HAM10000
QANA (Ours) 27.6 163.1 83.9 841.5 1.5 1.7 94.6 98.6
TABLE IV: Ablation study of core modules in our model on the HAM10000 test set. Metrics are reported as percentages (%). Each row shows the incremental addition of modules.
Configuration Ghost Block ECA SE Quant. Head SMOTE Inc. Learn Accuracy Recall Precision F1 Score AUC-ROC
@@
+ Incremental Learning 91.6 90.7 91.2 90.9 93.4
IV-E Inference Speed and Energy Consumption
We quantitatively evaluated the inference latency and energy consumption of our neuromorphic model in comparison with both conventional CNN baselines and other SNN-converted architectures. All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on both NVIDIA RTX 3090 GPU and an Intel Xeon CPU. For each model, the reported inference latency corresponds to the average per-image processing time over 1000 test samples. Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
To ensure a fair and hardware-consistent comparison, all tested CNN architectures were converted to SNNs using our pipeline before deployment on Akida. Table III presents a comprehensive summary of inference latency and energy consumption across all evaluated models and platforms, as well as the relative reduction of these metrics for SNNs on Akida compared to their CNN GPU and CPU implementations. Our neuromorphic model achieves the lowest inference latency and energy consumption of all evaluated architectures. When deployed as an SNN on the Akida platform, it completes classification in 1.5 ms per image and requires just 1.7 mJ, representing an 94.6% reduction in latency and over 98.6% reduction in energy compared to the equivalent CNN on GPU surpassing all other state-of-the-art CNN-to-SNN conversion baselines.
For images requiring higher spatial detail, the $\times2$ tiling fallback increases latency and energy approximately in proportion to the number of patches while remaining within the power envelope of the edge platform. This fallback is applied conditionally based on a confidence threshold on the single-pass prediction and does not change the training protocol.
TABLE V: Performance comparison of converted SNN models on the Clinical Dataset.

### v3

@@ 3.2.1 Iterative Feature Extraction and Downsampling
3.2.2 Spike-Compatible Feature Transformation
The backbone output $F^{(4)}$ is passed to a spike-compatible transformation module designed for direct quantization and SNN integration. We first apply a SeparableConv2D $(3\times 3,256)$ [31] to increase channel capacity while preserving hardware efficiency, then apply BatchNorm and an activation constraint that clips values into $[0,1]$:
$\widehat{F}=\min\left(1,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (7)
This bounded mapping aligns the activation range with spike encoding and avoids saturation under 8-bit affine quantization used in the subsequent conversion pipeline.
To further recalibrate channel responses, we apply a Squeeze-and-Excitation (SE) block [12] as shown in Fig. 5. The SE module computes channel-wise gates from global average pooled statistics using a two-layer bottleneck with a reduction ratio of $16$, then rescales $\widehat{F}$ by the resulting sigmoid gates. This improves class sensitivity under imbalance while keeping the operator set compatible with the Akida conversion and lowering flow.
3.2.3 Quantized Output Projection
Finally, the spike-compatible tensor is flattened to a vector $r\in\mathbb{R}^{4096}$ and projected to class logits through a linear readout:
$\mathbf{y}=\mathbf{W}_{\text{cls}}\,r+\mathbf{b}_{\text{cls}},$ (8)
where $\mathbf{W}_{\text{cls}}\in\mathbb{R}^{7\times 4096}$ and $\mathbf{b}_{\text{cls}}\in\mathbb{R}^{7}$. This projection is quantization-aware during training and produces logits in a form that is directly consumable by the CNN-to-SNN converter for neuromorphic inference.
3.3 CNN-to-SNN Conversion for Neuromorphic Deployment
We convert the trained quantization-aware CNN to an event-driven SNN with a graph-preserving flow that enforces an all-integer executable graph on Akida. The conversion consists of (i) static reparameterization, (ii) 8-bit affine quantization, and (iii) deterministic operator lowering.
BN folding and activation constraint. All BatchNorm layers are folded into the preceding convolution or pointwise projection. For output channel $c$ with convolution weights $\mathbf{W}_{c}$ and bias $b_{c}$ and BN parameters $(\gamma_{c},\beta_{c},\mu_{c},\sigma_{c}^{2},\epsilon)$, we use
@@ 3.3 CNN-to-SNN Conversion for Neuromorphic Deployment
Depthwise convolutions are folded channel-wise in the same manner. During training, non-supported nonlinearities are replaced by bounded ReLU6 so that post-activation ranges are fixed and compatible with spike encoding [15, 32].
@@
8-bit affine quantization. Weights and activations are quantized with uniform affine 8-bit fixed-point. For a real activation $a$ and calibration-derived scale and zero-point $(s_{a},z_{a})$, the quantized value is
@@
$a_{q}=\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{a}{s_{a}}\right)+z_{a},\,0,\,255\right),\qquad a\approx\left(a_{q}-z_{a}\right)s_{a}.$ (10)
For ReLU6-bounded activations, we set $z_{a}=0$ and $s_{a}=6/255$, which maps $[0,6]$ to $[0,255]$ without saturation. For layers constrained to $[0,1]$, including Eq. 7, we set $z_{a}=0$ and $s_{a}=1/255$.
Spike emission. SNN inference uses an integrate-and-fire neuron with subtractive reset driven by integer activations. With membrane state $u_{t}$, integer drive $a_{q}$, and threshold $\tau$, spike generation follows
$s_{t}=\mathbb{I}\!\left(u_{t-1}+a_{q}\geq\tau\right),\qquad u_{t}=\left(u_{t-1}+a_{q}\right)-\tau s_{t}.$ (11)
Class scores are obtained from spike counts over the integration window, and Softmax is applied after spike integration [32].
Integer graph lowering and metrics. Supported operators are lowered deterministically to integer kernels, including convolution (standard, depthwise, separable), pooling, elementwise clipping, residual addition with scale alignment, and linear layers. For residual addition, the two integer tensors are brought to a common affine scale before summation. Given $(s_{1},z_{1})$ and $(s_{2},z_{2})$, we choose a target scale $s_{\oplus}=\max(s_{1},s_{2})$ and rescale each branch by an integer multiplier and right shift so that both branches represent values under $s_{\oplus}$ prior to addition. This prevents overflow and reduces drift introduced by mixed-scale accumulation. Spatial reductions are compiled as integer reductions with the division absorbed into output scaling, and reshape operations do not introduce compute. Any subgraph that requires floating-point dequantization is rejected. We report conversion success as $r_{\mathrm{succ}}=1-\lvert\mathcal{V}_{\mathrm{fp32}}\rvert/\lvert\mathcal{V}\rvert$, and quantify conversion-induced changes by $\Delta\mathrm{Acc}=\mathrm{Acc}_{\mathrm{SNN}}-\mathrm{Acc}_{\mathrm{CNN}}$ and $\Delta\mathrm{F1}=\mathrm{F1}_{\mathrm{SNN}}-\mathrm{F1}_{\mathrm{CNN}}$ on identical splits and preprocessing.
3.4 SNN Deployment, Optimization and Inference
@@
3.4.1 Inference and Output Processing
@@
In our implementation, we use uniform temporal weights $w_{t}=1$ unless validation indicates a benefit from decaying integration. The scaling factor $\alpha$ is selected on the validation split and then fixed for the reported test evaluation.
3.4.2 Parameter Calibration and On-Chip Optimization
After initial deployment, key runtime parameters, including output spike thresholds, integration windows, and temporal pooling parameters, are empirically calibrated on a held-out validation set. To further optimize the class assignment and suppress spurious events, a threshold adaptation can be formulated as:
@@ 3.4.2 Parameter Calibration and On-Chip Optimization
where $S_{c}^{(i)}$ is the total spike count for class $c$ on sample $i$, $y_{i}$ is the true label, and $\mathbb{I}[\cdot]$ is the indicator function. This allows data-driven threshold selection for each class. If performance deviation is observed relative to the original CNN, light on-chip fine-tuning is conducted using Akida’s incremental learning capability, adjusting only the last output layer to optimize for domain shift or quantization artifacts. For higher-detail cases, a $2\times 2$ tiling fallback is triggered when $\max_{c}\hat{p}_{c}<0.60$. Four overlapping $64\times 64$ patches (20% stride overlap) are evaluated and logits averaged. On HAM10000, the trigger rate is 18%. All calibration, threshold selection, and optional on-chip fine-tuning steps use only the validation split. The test split is used only for final reporting and is not involved in any parameter selection.
4 Experiment
@@ 4.1 Datasets
We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from Nanjing Drum Tower Hospital, China. Both sets reflect a range of lesion types and clinical diversity.
4.1.1 HAM10000 Dataset
@@
The HAM10000 dataset [38] consists of 10,015 dermatoscopic RGB images labeled by expert dermatologists into 7 categories: melanocytic nevus, melanoma, benign keratosis-like lesions, basal cell carcinoma, actinic keratosis carcinoma, vascular lesions, and dermatofibroma. The images originate from diverse sources and exhibit significant class imbalance. For all experiments, we adopted a standard 70%/10%/20% train/validation/test split.
4.1.2 Hospital Clinical Dataset
A proprietary clinical dataset was established through a research collaboration with Nanjing Drum Tower Hospital, comprising 3,162 dermatoscopic images from 1,235 patients collected between June 2022 and February 2024 (Project No. NJG-2022-3233-CN). To ensure comparability with benchmark studies, lesion categories were curated to match the seven classes in HAM10000. Rare and unclassified conditions were excluded to maintain consistent labeling across classes. Each case was independently annotated by at least 2 board-certified dermatologists, with histopathological confirmation available for a subset of cases. All images and associated metadata were de-identified in accordance with ethics and governance requirements. Due to patient privacy, the dataset is not publicly released. Dataset partitioning followed a 70%/10%/20% train/validation/test split stratified by disease category. Splits were constructed to avoid patient overlap across training, validation, and test sets.
@@ 4.1.2 Hospital Clinical Dataset
Table 1: Class-wise precision, recall, F1 score of QANA on the HAM10000 test set.
Class Precision Recall F1
Actinic keratoses 0.890 0.933 0.911
Basal cell carcinoma 0.890 0.901 0.896
Benign keratosis-like lesions 0.866 0.853 0.859
Dermatofibroma 0.925 0.976 0.950
Melanocytic nevi 0.887 0.817 0.851
Vascular lesions 0.949 0.966 0.957
Melanoma 0.956 0.933 0.944
Average 0.909 0.911 0.910
Table 2: Performance Comparison of Converted SNN Models on HAM10000
@@
Ours 91.6 91.0
4.2 Experimental Setup
All model training, validation, and CNN-to-SNN conversion were conducted on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and NVIDIA RTX 3090 GPU, running Ubuntu 22.04 LTS. The software environment included Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1. Neuromorphic inference was performed on a BrainChip Akida AKD1000 PCIe board installed in the same system, with deployment and testing managed via the Akida Python API and default board settings.
4.3 Analysis of Classification Results on HAM10000 Dataset
Table 1 reports the class-wise precision, recall, and F1 score of QANA on the HAM10000 test set. Macro F1 is computed as the unweighted mean of per-class F1 scores on the test set. The model achieves consistent performance across lesion categories, including minority classes. With a Top-1 accuracy of 91.6% and a macro F1 of 91.0%, the system demonstrates effective discrimination of both common and rare lesions.
Table 3: Per-image inference latency and energy consumption of all models on the HAM10000 test set. CNN baselines are measured on an NVIDIA RTX 3090 (GPU) and Intel i9-12900K (CPU); SNNs are measured on BrainChip Akida AKD1000. All values are averaged over 10,000 images. The rightmost columns show the percentage reduction achieved by SNNs on Akida compared to the corresponding CNN version.
@@ 4.3 Analysis of Classification Results on HAM10000 Dataset
QANA (Ours) 27.6 163.1 83.9 841.5 1.5 1.7 94.6 99.0
Table 4: Ablation study of core modules in our model on the HAM10000 test set. Metrics are reported as percentages (%). Each row shows the incremental addition of modules. Inc. denotes fine-tuning of the readout layer after conversion using only validation data, while evaluation is performed on the fixed test split.
Configuration Ghost ECA SE Quant. SMOTE Inc. Acc. Rec. Prec. F1 AUC
@@
+ Incremental Learning 91.6 90.7 91.2 91.0 93.4
4.4 Classification Performance on HAM10000
To evaluate neuromorphic skin lesion classification under consistent deployment constraints, we trained a set of widely used CNN backbones and converted each model to an SNN using the same quantization-aware pipeline, then executed all converted networks on the same Akida platform with identical preprocessing and runtime settings. Table 2 summarizes the resulting Akida-SNN performance across canonical and advanced baselines, which typically exhibit reduced accuracy and macro F1 after conversion, with Top-1 accuracy spanning 83.2% to 88.1%. QANA attains 91.6% Top-1 accuracy and 91.0% macro F1 on HAM10000, yielding the best overall results under the same neuromorphic execution setting.
Table 5: Performance comparison of converted SNN models on the Clinical Dataset.